OpenStudy (anonymous):

Write the standard form of the equation of the circle with the given characteristics. Center: (0, 0); radius: 5

OpenStudy (anonymous):

Use \[ (x-a)^2 + (y-b)^2 =r^2 \]

OpenStudy (anonymous):

\[ x^2+y^2 =25 \]